OpenStudy (anonymous):

solving systems using substitution: whats x and y coordinates? -2x+5y=19 3x-4=y

OpenStudy (aripotta):

the second equation tells us what y equals in terms of x. so we can substitute that for y in the first equation -2x + 5y = 19 -2x + 5(3x - 4) = 19 now you can solve for x

OpenStudy (aripotta):

@isselpedroza you there?